OpenStudy (redneck88):

Can someone please explain how to figure out points of discontinuity? My book isnt doing me any good and i tried to look it up because its on my quiz but couldnt find any useful videos. Here's the question: Find any points of discontinuity for the rational function. y=(x+6)(x+2)(x+8) -------------- (x+9)(x+7)

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

set the denominator equal to zero and solve for \(x\) because you cannot divide by zero

OpenStudy (misty1212):

same question as "what is the domain of the expression"

OpenStudy (redneck88):

so how would i work that? x+9=0 and x+7=0?

OpenStudy (redneck88):

@misty1212

jimthompson5910 (jim_thompson5910):

from here, you solve x+9=0 for x to get ??? that's one x value that will make the denominator zero

jimthompson5910 (jim_thompson5910):

to get the other x value, you solve x+7=0 for x

OpenStudy (redneck88):

ok. so x=-9 and x=-7?

jimthompson5910 (jim_thompson5910):

correct, those two x values will make the entire denominator zero

jimthompson5910 (jim_thompson5910):

since you cannot divide by zero, they are the values that must be excluded from the domain (there are discontinuities here)

OpenStudy (redneck88):

ok.. does this make those two numbers points of discontinuity or is there more work to be done?

jimthompson5910 (jim_thompson5910):

there are no other points of discontinuity because you can plug in any other x value (other than x = -9 or x = -7) and you wouldn't have a division by zero error

OpenStudy (redneck88):

ah ok. thank you very much :)

jimthompson5910 (jim_thompson5910):

you're welcome